OpenStudy (anonymous):

2x^2 + 2y^2 + 2x - 4y = -1 find the standard form of the circle

OpenStudy (anonymous):

Do you know how to use completing the square? Because this is what you have to do in this case.

OpenStudy (anonymous):

Oh and divide by two first of course. makes it a bit easier.

OpenStudy (anonymous):

he's right, should be no problem solving now

OpenStudy (anonymous):

\[ k: f(x,y)=(x+\frac{1}{2})^2+(y-1)^2= \frac{3}{4} \] possible that I have done a careless mistake somewhere, but this is what I get.

OpenStudy (anonymous):

thanks